What is the height of a rectangular prism when the dimensions are:
-surface area = 438 inches squared
-the length of the base = 11 inches
-the width of base = 9 inches

Answer by venugopalramana(3286)

What is the height (H) of a rectangular prism when the dimensions are:
-surface area = 438 inches squared=A
-the length of the base = 11 inches=L
-the width of base = 9 inches=B
SURFACE AREA OF A RECTANGULAR PRISM=
A = 2(LB+BH+HL)=2(11*9+11H+9H)=2(99+20H)=198+40H
198+40H=438
40H=438-198=240
H=240/40=6